What would be the present value of an ending cash flow for a capital asset with an estimated salvage value of $30,000, which required $5,000 to be tied up in increased inventory levels during its life, if discounted at 7%, and the asset was expected to be used for another ten years?
Answer #1

Answer:

Project life is 10 years.

Ending (Terminal) cash flow = Salvage value + Recovery of working capital = 30000 + 5000 = $35,000

Ending cash flow correspond to Year 10

Discount rate given = 7%

Hence:

Present value of an ending cash flow = 35000 / (1 + 7%)^10 = $17,792.23

Present value of an ending cash flow = $17,792.23